When 25.0 mL of a 5.41×10-4 M manganese(II) acetate solution is combined with 12.0 mL of a 7.59×10-4 M potassium hydroxide solution does a precipitate form? (yes or no)

For these conditions the Reaction Quotient, Q, is equal to----?
Answer #1

It has the precipitate formation reaction:

Mn + 2 + 2 OH- = Mn (OH) 2

The concentration of reactive ions in solution is calculated:

[Mn + 2] = 5.41x10 ^ -4 M * 25 mL / 37 mL = 3.66x10 ^ -4 M

[OH-] = 7.59x10 ^ -4 M * 12 mL / 37 mL = 2.46x10 ^ -4 M

The quotient Q is calculated:

Q = [Mn + 2] * [OH -] ^ 2 = (3.66x10 ^ -4) * (2.46x10 ^ -4) ^ 2 = 2.21x10 ^ -11

You have to Qsp> Ksp, therefore if precipitate occurs.

If you liked the answer, please rate it positively, you would help me a lot, thanks.